how do you solve problem(8-20x)cubed=-1000?
Answer by checkley77(12844)

(8-20x)^3=-1000
taking the cube root of both sides we have.
8-20x=-10
-20x=-10-8
-20x=-18
x=-18/-20
x=9/10 answer.